OpenStudy (anonymous):

Complete the following: (a) Use the vertex formula to find the vertex of the graph of f(x)=1.5-4x-12x^2. (b) Find the intervals where f is increasing and where f is decreasing.

OpenStudy (anonymous):

find vertex via \[=\frac{b}{2a}=-\frac{-4}{2\times (-12)}=-\frac{1}{6}\]

OpenStudy (anonymous):

if you want the second coordinate of the vertex, which you do, you find \[f(\frac{1}{6})\]

OpenStudy (anonymous):

and since this is a parabola that opens down, it will be increasing on the interval \[(-\infty, \frac{1}{6})\] and decreasing on \[(\frac{1}{6},\infty)\]